A mutual fund company offers its customers a variety of funds: a money-market fund, three different bond funds (short, intermediate, and long-term), two stock funds (moderate and high-risk), and a balanced fund. Among customers who own shares in just one fund, the percentages of customers in the different funds are as follows.

Money-market 22%    High-risk stock 16%
Short bond 12%    Moderate-risk stock   25%
Intermediate bond   5%    Balanced 15%
Long bond 5%               

A customer who owns shares in just one fund is randomly selected.

(a) What is the probability that the selected individual owns shares in the balanced fund?


(b) What is the probability that the individual owns shares in a bond fund?


(c) What is the probability that the selected individual does not own shares in a stock fund?
Answer #1

A) P(balanced fund) = 0.15

B) P(bond fund) = P(short bond) + P(intermediate bond) + P(long bond)

= 0.12 + 0.05 + 0.05

= 0.22

C) P(not stock fund) = 1 - P(stock fund)

= 1 - (P(high risk stock) + P(moderate risk stock))

= 1 - (0.16 + 0.25)

= 0.59